OpenStudy (anonymous):

in a farm yard there are only chickens and sheep. there are 13 animals altogether and between them they have 34 legs. use simultaneous equation to determine how many chickens and sheep there are.

OpenStudy (anonymous):

x chiken y sheep \[\begin{cases}x+y=13\\2x+4y=34\end{cases}\]

OpenStudy (anonymous):

since chicken has 2 legs and sheep has 4